OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 5√8 inches and a width of 2√5 inches.

OpenStudy (anonymous):

A =lw A=(5√8)( 2√5) A=10√40 A=10√4*10 A=10*2√10 A=20√10